# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v2

@@ 2.2 Energy model
Figure 2: The accuracy of the energy model and the calculated specific heats for FeCoNiAlTi and MoNbTaW. (a) The training error of the EPI model for FeCoNiAlTi. (b) The validation error of the EPI model for FeCoNiAlTi (0.7/0.3 train-valid splitting). (c) The training error of the loca_SRO model for FeCoNiAlTi. (d) The validation error of the local_SRO model for FeCoNiAlTi. (e) The specific heats $C_{V}$ of FeCoNiAlTi were calculated with different supercell sizes. (f) The specific heats of MoNbTaW were calculated with a one-million-atom supercell.
@@
Using a train-validation splitting of 70% and 30% for the DFT dataset described in Method. 3.4, we evaluated the accuracy of the second-nearest-neighbor (2nd-NN) EPI model and the loca_SRO model and the results are shown in Fig. 2 (a-d). Both models are trained with the Adam optimizer and back-propagation algorithm. It can be seen that both of the two models demonstrate very high accuracy. For the EPI model, the validation root-mean-square-error (RMSE) is 0.1819 mRy, or approximately 2.5 meV, which is well within quantum chemical accuracy (approximately 3.16 mRy/atom), as well as smaller than the typical errors of $5-12$ meV/atom in the DFT methods [60]. The local_SRO model demonstrates slightly higher accuracy than the EPI model, with a validation RMSE of 0.1643 mRy/atom, which can be attributed to its inclusion of high-order interactions. The high accuracy of the models is also demonstrated in the $R^{2}$ scores, which are all higher than 0.995. The errors from the train and validation datasets are also very close to each other, indicating the absence of overfitting in the models. For comparison, we notice that this accuracy is higher than that in Ref. [29], which is reported as 14.7 meV/atom by applying a similar pair-interaction model in the FeCoNiAlTiB system. We think one reason for the enhanced accuracy is our approach to obtaining the DFT dataset, which combines both random configurations, as well as configurations from Monte Carlo simulations, as proposed in Ref. [33] and described in Method. 3.4.
@@ 2.3 Nanostructures in $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$
In this section, we report the nanostructure evolution of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc HEA. This alloy has demonstrated exceptional combined strength and ductility in experiments, and such an attractive property has been attributed to the formation of nanoparticles in the FeCoNi matrix [39], which hinder the motion of dislocations. In the experiment, the alloy sample is homogenized at 1150 ${}^{\circ}C$ (1423 K) for 2 hours and aged at 780 ${}^{\circ}C$ (1053 K) for 4 hours [39]. Using the local SRO model, we first simulated the chemical phase changes of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc alloy using a supercell of approximately one million atoms. The lattice dimension is chosen as $72\times 60\times 60$, therefore, the total number of atoms is $72\times 60\times 60\times 4=1,036,800$. We initialize the configurations randomly at 2000 K, then decrease the simulation temperature by 50 K each time until 1000 K, which is approximately the temperature at which the samples are aged. At each temperature, we skip the $4\times 10^{5}-1.5\times 10^{6}$ MC sweeps before recording the following 20,000 sweeps to calculate the specific heat $C_{V}$ in Fig. 2 (e). The configurations at 2000 K and 1000 K are demonstrated in Fig. 3 (a-e). From Fig. 3 (d-e) it can be seen that the system is chemically disordered at 2000 K. As the temperature decreases to 1000 K, a nanoparticle (NP) forms in the system. The distribution of elements in the NP, as shown in Fig. 3 (a-b), reveals the formation of $L1_{2}$ structure. The two different layers of the $L1_{2}$ nanoparticle can be seen from Fig. 3 (b-c). This agrees well with the experimental results in Ref. [39], as reproduced in Fig. 3 (f-g), in which the nanoparticles are refered to as multi-component intermetallic nanoparticles (MCINP). Similar results are also reported in Ref. [63] for the $\rm{Fe_{25}Co_{25}Ni_{25}Al_{15}Ti_{10}}$ HEA, where the fcc phase consisted of a $\gamma$ Fe-(Co,Ni)-based solid-solution matrix (A1), and coherent primary $\gamma^{\prime}$ (Ni,Co)3-(Ti,Al)-based intermetallic $L1_{2}$ precipitates. Other than chemical structure, size is another important feature of the nanoparticles. The diameter of the cluster shown in Fig. 3 (a) is about 17.8 nm, which is in agreement with the experimental observation shown in Fig. 3 (f-g). Moreover, we extract the compositions from the simulation data and show it in Fig. 4 (g), along with the experiment results in Ref. [39], as shown in Fig. 4 (h). It can be seen that the compositions from MC simulation generally agrees well with experiment, although there are also some differences, such as slightly higher concentration of Ni in MCINP.
@@
Figure 4: Comparing the simulated configurations of the 1 billion atom FeCoNiAlTi with the experimental results. (a) A histogram for all clusters of sizes larger than 100 atoms. Five clusters of different sizes are selected and shown as vertical dashed lines. (b) The chemical concentrations of the matrix (cluster_0) and cluster_1, as well as the rest clusters. (c-f) Snapshots of the [001] face of NiCoFeAlTi at T=1000K, with (c) as the enlarged upper-right corner of (d), and (e) as the enlarged upper-right corner of (f). (c-d) show the different elements, and (e-f) show the atomic local energies. (g) Compositions of the NP (MCINP) and matrix (MCM) from the 1M atom simulation via sampling along the x direction using a radius of 6.4 nm. (h) Compositions of the NP (MCINP) and matrix (MCM) from experiment, as reproduced from Ref. [39].
@@
Figure 5: Comparison between the NPs from MC simulation and APT experiment. (a-c) 3D snapshot of cluster_4, 3, and 2, respectively. (d) APT results from fcc grains of FNAT sample (47-hours aging at 973 K). (e) Same as (d), but for a sample obtained via 4-hours aging. Both figure (d) and (e) are removed and shown as blank on purpose for permission reasons in this preprint. See original ones at Fig. 2 of Ref. [40]. (f) The APT results of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA, as reproduced from [39]. (g) The simulation result obtained using SMC. The blue region show the NP identified, and the red region show the matrix phase. (h) The simulation result obtained with SMC. The chemical species are explicitly shown.
@@
From Fig. 3 (b,c), we see that the one-million-atom supercell can only accommodate one NP in the FeCoNiAlTi HEA, and larger systems are necessary for a direct comparison of experimental data that contains more than one NP, such as the TEM image in Fig. 3 (g). To investigate the microstructures of the FeCoNiAlTi HEA at a larger length scale, we employ a $630\times 630\times 630$ fcc supercell, in which the total number of atoms is 1,000,188,000. Due to the huge supercell size, we use the faster EPI model rather than the local SRO model, and the simulation details are specified in Method. 3.5. The simulation results for this one billion-atom system are presented in Fig. 4 and Fig. 5. To study the size, composition, and morphologies of the nanoparticles, we employ the union-find algorithm to identify the nanoparticles, as detailed in Method. 3.5. We refer to all sets of atoms identified by the union-find algorithm as clusters, which can be of sizes ranging from a single isolated atom to the whole matrix phase. We make a histogram of all clusters of sizes larger than 100 atoms in Fig. 4 (a). The largest one is cluster_0, which is the matrix phase. Cluster_0 has a total of 0.6 billion atoms, and contains only Fe, Co, and Ni, as demonstrated in Fig. 4 (b). Note that the absence of Al and Ti atoms in cluster_0 is a result of the union-find algorithm, which automatically identifies isolated Al and Ti atoms in cluster_0 as individual clusters of size one. From Fig. 4 (c), we see that the FeCoNi matrix phase is a disordered A1 structure. The cause of the disorder can be seen from Fig. 4 (e) and (f), which show the local energies of each atom with the color bar. It can be seen that in the matrix phase, the local energies of different atoms are generally close to each other, which means that the Fe, Co, and Ni atoms have no significant site preferences. On the other hand, the nanoparticles, which forms $L1_{2}$ structure as shown in Fig. 4 (c), contains sites with much lower atomic local energies, as shown in Fig. 4 (e). In fact, the local energy difference is exactly the criteria we make use of to distinguish the matrix and nanoparticles, as described in Method 3.5. The FeCoNi matrix phase is the entropy-stablized disordered phase, while the $L1_{2}$ phase is the enthalpy favorable ordered phase. In other words, the occurrence of both disordered A1 structure and ordered $L1_{2}$ structure is a result of the competition between the entropy and enthalpy contributions in the free energy.
@@
To study the size, compositions, and morphologies of the nanoparticle in more detail, we choose four representative nanoparticles, as illustrated in Fig. 4 (a) as cluster_1, 2, 3, and 4, which have 264,792,042, 1,007,916, 102,389, and 20,937 atoms. Their chemical concentrations are also listed in Tab. 2, which aligns well with the experimental results [29]. From Fig. 4 (a), we note that the peak of the histogram is at about $10^{5}$ atoms, which corresponds to a cluster of 29 atoms, or 10.3 nm in diameter. For cluster_1, a similar calculation gives a value of 22.0 nm in diameter, if we assume that the particles are spherical. However, inspecting the 3D morphology of the clusters reveals that they are actually not simply spheres or ellipsoids, as shown in Fig. 5 (a-c): While ellipsoid is a good approximation for the smaller cluster_4, the larger cluster_3 and cluster_2 are both comprising of multiple nanoparticles, which indicates that the smaller nanoparticles can merge together to form larger ones, as shown by the dendritic structures at the surface of the nanoparticles. In fact, from Fig. 4 (a,b) we see that the largest $L1_{2}$ cluster: cluster_1, contains a total of more than 264 million atoms, which accounts for 66% of the atoms in the NP phases. In other words, most of the nanoparticles shown in Fig. 4 (c-f) are actually connected at 3-dimension and are parts of cluster_1. Compared to simple NP, we propose that this 3D-connected-NPs (3DCNP) can further hinder the motion of dislocations, which can serve as a strengthening mechanism in these types of HEAs. Verifying the existence of the 3DCNP demands high-precision experimental techniques such as atom probe tomography (APT). The APT image of $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is available in Ref. [39], as reproduced in Fig. 5 (f). However, the image only shows the surface of the sample instead of the 3D chemical structure. Although the theoretical observation of the 3DCNP still requires experimental validation, we contend that its existence is plausible, given the high density of NPs observed in experiment [39]. Moreover, the APT result of the FNAT alloy [40], a medium entropy alloy made up of the Fe, Ni, Al, and Ti elements, seems to support the existence of $L1_{2}$ 3DCNP, as shown in Fig. 5 (d,e). For a direct comparison with the experimental results, we also show the simulation results of the APT sample needle in Fig. 5 (g, h). It is easy to see that the general shapes of the NPs from simulation resemble the experimental results of FNAT, despite their different chemical compositions. The sizes of the NPs from simulation is generally smaller than experimental results, which could be due to the relatively limited $10^{5}$ MC simulation steps. Again, the importance of SMC method is highlighted by the capability to directly simulate an APT needle comprising of 10 million atoms. Other than the mechanical properties, the NPs can also have important application in catalysis [64], which would be an interesting topic for future research [46, 65].
@@
[…]

### v3

@@ 2.2 Energy model
Figure 2: The accuracy of the energy model (RMSE per atom) and the calculated specific heats for FeCoNiAlTi and MoNbTaW. (a) The training error of the EPI model for FeCoNiAlTi. (b) The validation error of the EPI model for FeCoNiAlTi (0.7/0.3 train-valid splitting). (c) The training error of the loca_SRO model for FeCoNiAlTi. (d) The validation error of the qSRO model for FeCoNiAlTi. (e) The specific heats $C_{V}$ of FeCoNiAlTi were calculated with different supercell sizes. (f) The specific heats of MoNbTaW were calculated with a one-million-atom supercell.
@@
Using a train-validation splitting of 70% and 30% for the DFT dataset described in Method. 3.4, we evaluated the accuracy of the second-nearest-neighbor (2nd-NN) EPI model and the loca_SRO model and the results are shown in Fig. 2 (a-d). Both models are trained with the Adam optimizer and back-propagation algorithm. It can be seen that both of the two models demonstrate very high accuracy. For the EPI model, the validation root-mean-square-error (RMSE) is 0.1819 mRy/atom, or approximately 2.5 meV/atom, which is well within quantum chemical accuracy (approximately 3.16 mRy/atom), as well as smaller than the typical errors of 5125-12 meV/atom in the DFT methods [72]. The qSRO model demonstrates slightly higher accuracy than the EPI model, with a validation RMSE of 0.1643 mRy/atom, which can be attributed to its inclusion of high-order interactions. The high accuracy of the models is also demonstrated in the $R^{2}$ scores, which are all higher than 0.995. The errors from the train and validation datasets are also very close to each other, indicating the absence of overfitting in the models. For comparison, we notice that this accuracy is higher than that in Ref. [29], which is reported as 14.7 meV/atom by applying a similar pair-interaction model in the FeCoNiAlTiB system. We think one reason for the enhanced accuracy is our approach to obtaining the DFT dataset, which combines both random configurations, as well as configurations from Monte Carlo simulations, as proposed in Ref. [33] and described in Method. 3.4.
@@ 2.3 Nanostructures in $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$
In this section, we report the nanostructure evolution of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc HEA. This alloy has demonstrated exceptional combined strength and ductility in experiments, and such an attractive property has been attributed to the formation of nanoparticles in the FeCoNi matrix [54], which hinder the motion of dislocations. In the experiment, the alloy sample is homogenized at 1150 $^{\circ}C$ (1423 K) for 2 hours and aged at 780 $^{\circ}C$ (1053 K) for 4 hours [54]. Using the local SRO model, we first simulated the chemical phase changes of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc alloy using a supercell of approximately one million atoms. The lattice dimension is chosen as $72\times 60\times 60$, therefore, the total number of atoms is $72\times 60\times 60\times 4=1,036,800$. We initialize the configurations randomly at 2000 K, then decrease the simulation temperature by 50 K each time until 1000 K, which is approximately the temperature at which the samples are aged. At each temperature, we skip the $4\times 10^{5}-1.5\times 10^{6}$ MC sweeps before recording the following 20,000 sweeps to calculate the specific heat $C_{V}$ in Fig. 2 (e). The configurations at 2000 K and 1000 K are demonstrated in Fig. 3 (a-e). From Fig. 3 (d-e) it can be seen that the system is chemically disordered at 2000 K. As the temperature decreases to 1000 K, a nanoparticle (NP) forms in the system. The distribution of elements in the NP, as shown in Fig. 3 (a-b), reveals the formation of $\text{L1}_{2}$ structure. The two different layers of the $\text{L1}_{2}$ nanoparticle can be seen from Fig. 3 (b-c). This agrees well with the experimental results in Ref. [54], as reproduced in Fig. 3 (f-g), in which the nanoparticles are refered to as multi-component intermetallic nanoparticles (MCINP). Similar results are also reported in Ref. [75] for the $\rm{Fe_{25}Co_{25}Ni_{25}Al_{15}Ti_{10}}$ HEA, where the fcc phase consisted of a $\gamma$ Fe-(Co,Ni)-based solid-solution matrix (A1), and coherent primary $\gamma^{\prime}$ (Ni,Co)3-(Ti,Al)-based intermetallic $\text{L1}_{2}$ precipitates. Other than chemical structure, size is another important feature of the nanoparticles. The diameter of the cluster shown in Fig. 3 (a) is about 17.8 nm, which is in agreement with the experimental observation shown in Fig. 3 (f-g). Moreover, we extract the compositions from the simulation data and show it in Fig. 4 (g), along with the experiment results in Ref. [54], as shown in Fig. 4 (h). It can be seen that the compositions from MC simulation generally agrees well with experiment, although there are also some differences, such as slightly higher concentration of Ni in MCINP.
@@
Figure 4: Comparing the simulated configurations of the 1 billion atom FeCoNiAlTi with the experimental results. (a) A histogram for all clusters of sizes larger than 100 atoms. Five clusters of different sizes are selected and shown as vertical dashed lines. (b) The chemical concentrations of the matrix (cluster_0) and cluster_1, as well as the rest clusters. (c-f) Snapshots of the [001] face of NiCoFeAlTi at T=1000K, with (c) as the enlarged upper-right corner of (d), and (e) as the enlarged upper-right corner of (f). (c-d) show the different elements, and (e-f) show the atomic local energies. (g) Compositions of the NP (MCINP) and matrix (MCM) from the 1M atom simulation via sampling along the x direction using a radius of 6.4 nm. (h) Compositions of the NP (MCINP) and matrix (MCM) from experiment, as reproduced from Ref. [54] with permission from Science.
@@
Figure 5: Comparison between the NPs from MC simulation and APT experiment. (a-c) 3D snapshot of cluster_4, 3, and 2, respectively. (d) APT results from fcc grains of FNAT sample (47-hours aging at 973 K) showing the distribution and composition of $\text{L1}_{2}$ precipitates in fcc matrix. (e) Same as (d), but for a sample obtained via 4-hours aging. Both (d) and (e) are reproduced from Ref. [55]. We will seek permission once this manuscript being accepted. (f) The APT results of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA, as reproduced from [54]. (g) The simulation result obtained using SMC-X. The blue region show the NP identified, and the red region show the matrix phase. (h) The simulation result obtained with SMC-X. The chemical species are explicitly shown.
@@
From Fig. 3 (b,c), we see that the one-million-atom supercell can only accommodate one NP in the FeCoNiAlTi HEA, and larger systems are necessary for a direct comparison of experimental data that contains more than one NP, such as the TEM image in Fig. 3 (g). To investigate the microstructures of the FeCoNiAlTi HEA at a larger length scale, we employ a $630\times 630\times 630$ fcc supercell, in which the total number of atoms is 1,000,188,000. Due to the huge supercell size, we use the faster EPI model rather than the local SRO model, and the simulation details are specified in Method. 3.5. The simulation results for this one billion-atom system are presented in Fig. 4 and Fig. 5. To study the size, composition, and morphologies of the nanoparticles, we employ the union-find algorithm to identify the nanoparticles, as detailed in Method. 3.5. We refer to all sets of atoms identified by the union-find algorithm as clusters, which can be of sizes ranging from a single isolated atom to the whole matrix phase. We make a histogram of all clusters of sizes larger than 100 atoms in Fig. 4 (a). The largest one is cluster_0, which is the matrix phase. Cluster_0 has a total of 0.6 billion atoms, and contains only Fe, Co, and Ni, as demonstrated in Fig. 4 (b). Note that the absence of Al and Ti atoms in cluster_0 is a result of the union-find algorithm, which automatically identifies isolated Al and Ti atoms in cluster_0 as individual clusters of size one. From Fig. 4 (c), we see that the FeCoNi matrix phase is a disordered A1 structure. The cause of the disorder can be seen from Fig. 4 (e) and (f), which show the local energies of each atom with the color bar. It can be seen that in the matrix phase, the local energies of different atoms are generally close to each other, which means that the Fe, Co, and Ni atoms have no significant site preferences. On the other hand, the nanoparticles, which forms $\text{L1}_{2}$ structure as shown in Fig. 4 (c), contains sites with much lower atomic local energies, as shown in Fig. 4 (e). In fact, the local energy difference is exactly the criteria we make use of to distinguish the matrix and nanoparticles, as described in Method 3.5. The FeCoNi matrix phase is the entropy-stablized disordered phase, while the $\text{L1}_{2}$ phase is the enthalpy favorable ordered phase. In other words, the occurrence of both disordered A1 structure and ordered $\text{L1}_{2}$ structure is a result of the competition between the entropy and enthalpy contributions in the free energy.
@@
To study the size, compositions, and morphologies of the nanoparticle in more detail, we choose four representative nanoparticles, as illustrated in Fig. 4 (a) as cluster_1, 2, 3, and 4, which have 264,792,042, 1,007,916, 102,389, and 20,937 atoms. Their chemical concentrations are also listed in Tab. 3, which aligns well with the experimental results [29]. From Fig. 4 (a), we note that the peak of the histogram is at about $10^{5}$ atoms, which corresponds to a cluster of 29 atoms, or 10.3 nm in diameter. For cluster_1, a similar calculation gives a value of 22.0 nm in diameter, if we assume that the particles are spherical. However, inspecting the 3D morphology of the clusters reveals that they are actually not simply spheres or ellipsoids, as shown in Fig. 5 (a-c): While ellipsoid is a good approximation for the smaller cluster_4, the larger cluster_3 and cluster_2 are both comprising of multiple nanoparticles, which indicates that the smaller nanoparticles can merge together to form larger ones, as shown by the dendritic structures at the surface of the nanoparticles. In fact, from Fig. 4 (a,b) we see that the largest $\text{L1}_{2}$ cluster: cluster_1, contains a total of more than 264 million atoms, which accounts for 66% of the atoms in the NP phases. In other words, most of the nanoparticles shown in Fig. 4 (c-f) are actually connected at 3-dimension and are parts of cluster_1. Compared to simple NP, we propose that this 3D-connected-NPs (3DCNP) can further hinder the motion of dislocations, which can serve as a strengthening mechanism in these types of HEAs. Verifying the existence of the 3DCNP demands high-precision experimental techniques such as atom probe tomography (APT). The APT image of $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is available in Ref. [54], as reproduced in Fig. 5 (f). However, the image only shows the surface of the sample instead of the 3D chemical structure. Although the theoretical observation of the 3DCNP still requires experimental validation, we contend that its existence is plausible, given the high density of NPs observed in experiment [54]. Moreover, the APT result of the FNAT alloy [55], a medium entropy alloy made up of the Fe, Ni, Al, and Ti elements, seems to support the existence of $\text{L1}_{2}$ 3DCNP, as shown in Fig. 5 (d,e). For a direct comparison with the experimental results, we also show the simulation results of the APT sample needle in Fig. 5 (g, h). It is easy to see that the general shapes of the NPs from simulation resemble the experimental results of FNAT, despite their different chemical compositions. The sizes of the NPs from simulation is generally smaller than experimental results, which could be due to the relatively limited $10^{5}$ MC simulation steps. Again, the importance of SMC-X method is highlighted by the capability to directly simulate an APT needle comprising of 10 million atoms. Other than the mechanical properties, the NPs can also have important application in catalysis [76], which would be an interesting topic for future research [60, 77].
@@
Table 3: Comparison of the experimental values at 1055 K and theoretical values at 1100 K from Ref. [29] with our simulation results at 1000 K for the compositions of the matrix and four representative $\rm{\text{L1}_{2}}$ clusters, as listed in Fig. 4. Note that $\rm{Matrix}^{MC}$ is Cluster_0. Due to the adoption of the union-find algorithm, the isolated atoms are automatically removed from the identified clusters, which is the cause of no Al and Ti atoms in the extracted matrix phase.
@@ 2.4 Nanostructures in MoNbTaW
2.5 Discussion of limits
To facilitate better interpretation of the simulation results, we examine the method’s limitations and suggest directions for future refinement. Note that the interaction range used in this work is limited to the first and second nearest neighbors. In contrast, Ref. [78] suggests that effective pair interactions (EPIs) can be long-ranged, with contributions from up to 99 coordination shells ($p_{\text{max}}$) required for accurate total energy estimation. However, those EPIs were derived using the EMTO+CPA electronic structure method and are not necessarily equivalent to the EPIs used in this work, which were obtained by regression from direct DFT calculations. In a previous study, Liu et al. investigated the influence of interaction range on model performance. Their findings indicate that the optimal cutoff depends on the training dataset, governed by the well-known bias–variance trade-off. In general, the first two coordination shells capture the most significant contributions to the energy, while extending the range can improve accuracy if supported by a larger DFT dataset. Comparing this work, which uses a cutoff of $p_{\text{max}}=2$, to Ref. [78], which used $p_{\text{max}}=6$, the resulting order–disorder transition temperatures are 1050 K and 873 K, respectively. This relatively modest difference supports the adequacy of a short-range model for the current study, given the size of the available DFT dataset.
A further limitation of this work is that lattice distortions are not included in the DFT calculations, which can potentially change the order-disorder transition temperature significantly, as noticed for MoNbTaW [28]. The lack of lattice relaxation in our DFT data stems from the use of the LSMS method, which, while computationally efficient, does not currently support lattice relaxations. Although the distortions in the HEAs studied here are generally small, their precise impact should be addressed in future work.
Finally, beyond lattice distortion, several other factors can influence phase transitions. A rigorous treatment of these effects requires evaluating all relevant contributions to the Gibbs free energy, including electronic, vibrational (phonon), and configurational terms. When the crystal structure remains unchanged, it is generally reasonable to approximate phase stability using only the configurational term, as chemical disorder predominantly introduces smearing of the Fermi surface and phonon dispersion. Moreover, volume thermal expansion is typically less than 1% between room temperature and 1000 K [79], and such a minor change alone is unlikely to significantly affect the phase transition behavior. Nonetheless, a comprehensive evaluation of all relevant contributions is required for a definitive conclusion, which is beyond the scope of this study.